OpenStudy (anonymous):

Find the margin of error when p = 83% and n = 300. Round your answer to the nearest hundredth.

OpenStudy (anonymous):

Hmm, don't you need a confidence level to determine the margin of error?

OpenStudy (anonymous):

I think so

OpenStudy (anonymous):

Weird... well, I would assume a 95% confidence level (that's the standard one I think), which would make the margin of error \[1.96\sqrt\frac{p(1-p)}{ n}\] but that depends on the confidence level, like I mentioned.

OpenStudy (anonymous):

is 1.96 the answer?

OpenStudy (anonymous):

For the margin of error? No. The margin of error (*assuming* 95% confidence) would be \[1.96\sqrt\frac{.83(1-.83)}{300}\]

OpenStudy (anonymous):

can you calculate that

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):

thanks